# On the tails of log-concave density estimators

### v2

@@ Abstract.
It is shown that the nonparametric maximum likelihood estimator of a univariate log-concave probability density satisfies some consistency properties in the tail regions.
@@
Chernov bounds, consistency, exponential inequalties, log-linear densities.
@@ 1 Introduction
Concerning consistency of $\hat{\varphi}_{n}$, let $\{x\in\mathbb{R}:0<F(x)<1\}=:(a_{o},b_{o})$ with $\infty\leq a_{o}<b_{o}\leq\infty$. It was shown by Dümbgen and Rufibach (2009) that for any fixed interval $a,b]\subset(a_{o},b_{o})$, the supremum of $|\hat{\varphi}_{n}-\varphi|$ over $[a,b]$ is of order $O_{\mathrm{p}}(\rho_{n}^{1/3})$, where $\rho_{n}:=\log(n)/n$. If $\varphi$ is Hölder-continuous on a neighborhood of $[a,b]$ with exponent $\beta\in(1,2]$, this rate improves to $O_{\mathrm{p}}(\rho_{n}^{\beta/(2\beta+1)})$. Uniform consistency of $\hat{\varphi}_{n}$ on arbitrary compact subintervals of $(a_{o},b_{o})$ implies that $\int\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\,\mathrm{d}x\to_{\mathrm{p}}0$. (Throughout this note, asymptotic statements refer to $n\to\infty$.) Pointwise limiting distributions at a single point $x_{o}\in(a_{o},b_{o})$ have been derived by Balabdaoui et al. (2009), assuming that $\varphi$ is twice continuously differentiable in a neighborhood of $x_{o}$. Kim and Samworth (2016) showed that the expected squared Hellinger distance between $\hat{f}_{n}$ and $f$ is of order $O(n^{-4/5})$. Numerous further results about $\hat{\varphi}_{n}$ and $\hat{f}_{n}$ have been derived thereafter, including multivariate settings, see the review of Samworth (2018). In the present univariate setting, fast algorithms for the computation of $\hat{\varphi}_{n}$ and related objects are provided by Dümbgen and Rufibach (2011) and Dümbgen et al. (2021). Experiments with simulated data show that even in the tail regions, that is, close to $a_{o}$ and $b_{o}$, the estimator $\hat{\varphi}_{n}$ is surprisingly accurate. In view of this empirical finding, Müller and Rufibach (2009) developed new estimators for extreme value analysis with excellent empirical performance. However, the currently available theory about the asymptotic properties of $\hat{\varphi}_{n}$ does not explain its good performance in the tail regions. The present note provides a first step into that direction by proving several results about $\hat{\varphi}_{n}(x)$ and the right-sided derivative $\hat{\varphi}_{n}^{\prime}(x+)$ for $x$ close to $b_{o}$. By symmetry, these findings carry over to results in the left tail region. The main results are presented in Section 2 while Sections 3 and 4 provide the proofs.
@@ 2 Main results
(c)  Suppose that $\varphi^{\prime}(\infty-)=$-$\infty$ and $\varphi$ is differentiable on some halfline $(a_{*},\infty)\subset(a_{o},\infty)$ with Lipschitz-continuous derivative $\varphi^{\prime}$. Then, for arbitary fixed $a\in(a_{*},\infty)$ such that $\varphi^{\prime}(a)<0$,
$\sup_{x\in[a,b_{n}]}\,\Bigl{(}\frac{\hat{\varphi}_{n}^{\prime}(x+)}{\varphi^{\prime}(x)}-1\Bigl{)}^{+}\$$\to_{\mathrm{p}}\ 0,$
$\sup_{x\in[a,b_{n}]}\,\frac{\bigl{(}\varphi(x)-\hat{\varphi}_{n}(x)\bigr{)}^{+}}{1+|\varphi(x)|}\$$\to_{\mathrm{p}}\ 0,$
$\sup_{x\in[a,b_{n}]}\,\bigl{(}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{)}^{+}\$$\to_{\mathrm{p}}\ 0.$
@@ 3.2 Inequalities for $\hat{P}_{n}^{\mathrm{emp}}$
Note first that for arbitrary indices $0\leq j<k\leq n+1$ with $k-j\leq n$, the random variable $P(X_{(j)},X_{(k)})$ follows a beta distribution with parameters kjk-j and n+1k+jn+1-k+j, see Chapter 3.1 of Shorack and Wellner (1986). In particular, its mean equals
@@ 3.3 Truncated and conditional means of $P$
Note that $x_{1}=a$ would imply that $f>\tilde{f}$ on $(a,b)$ or $f<\tilde{f}$ on $(b,\infty)$, and in both cases we would end up with $\int_{a}^{b}f(x)\,\mathrm{d}x\neq\int_{a}^{b}\tilde{f}(x)\,\mathrm{d}x$. Similiarly one can exclude the cases $x_{2}=\infty$, $x_{1}\geq b$ and $x_{2}\leq b$. Consequently, we know that there exist constants $a<x_{1}<b<x_{2}<\infty$ such that
@@ 3.4 Exponential and maximal inequalities
If $Q$ follows an exponential distribution, it is well known that the above inequality holds with equality.
Now suppose that $Q$ is an arbitrary distribution with log-density $\varphi$. Let $\tilde{\varphi}$ be the log-density of the exponential distribution $\tilde{Q}$ with mean $\mu_{Q}$, that is, $\tilde{\varphi}(x)=-x/\mu_{Q}-\log(\mu_{Q})$ for $x\geq 0$ and $\tilde{\varphi}(x)=$-$\infty$ for $x<0$. If $\varphi\not\equiv\tilde{\varphi}$, then there exist points $0<a<b<\infty$ such that $\varphi<\tilde{\varphi}$ on $\mathbb{R}\setminus[a,b]$ and $\varphi>\tilde{\varphi}$ on $(a,b)$. Indeed, by concavity of $\varphi-\tilde{\varphi}$ on $0,\infty)$, the set $\{\varphi>\tilde{\varphi}\}$ is an interval with infimum $a\geq 0$ and supremum $b\leq\infty$, and $\varphi<\tilde{\varphi}$ on $0,\infty)\setminus[a,b]$. Since $\int e^{\varphi(x)}\,\mathrm{d}x=\int e^{\tilde{\varphi}(x)}\,\mathrm{d}x=1$, we know that $a<b$ and $(a,b)\neq(0,\infty)$. If $a=0$, then the distribution function $F$ of $Q$ would be strictly larger than the distribution function $\tilde{F}$ of $\tilde{Q}$ on $(0,\infty)$, leading to the contradiction that $\mu_{Q}=\int_{0}^{\infty}(1-F(x))\,\mathrm{d}x<\mu_{Q}$. With the same argument one can argue that $b<\infty$. But now, for any convex function $\Psi$ on the real line,
$\int\Psi\,\mathrm{d}Q\ \leq\ \int\Psi\,\mathrm{d}\tilde{Q}$. (11)
Indeed, since $\int(c+dx)\,P(\mathrm{d}x)=\int(c+dx)\,\tilde{P}(\mathrm{d}x)$ for arbitrary $c,d\in\mathbb{R}$, the difference $\int\Psi\,dQ-\int\Psi\,\mathrm{d}\tilde{Q}$ equals
$\int\Psi(x)(e^{\varphi(x)}-e^{\tilde{\varphi}(x)})\,\mathrm{d}x\ =\ \int\Bigl{(}\Psi(x)-\frac{b-x}{b-a}\Psi(a)-\frac{x-a}{b-a}\Psi(b)\Bigr{)}(e^{\varphi(x)}-e^{\tilde{\varphi}(x)})\,\mathrm{d}x\ \leq\ 0,$
because the integrand of the latter integral is non-positive. Applying (11) to $\Psi(x)=\exp(tx)$ for arbitrary $t<1$, and using independence of $Y_{1},\ldots,Y_{m}$ concludes the proof. ∎
@@ 4 Proofs of the main results
To this end, we show that for any fixed $\varepsilon>0$, the inequality $\hat{\varphi}_{n}(b_{n})\leq\varphi(b_{o})-\varepsilon$ holds with asymptotic probability zero. Let $b(\varepsilon)\in(a_{o},b_{o})$ such that $|\varphi-\varphi(b_{o})|\leq\lambda\varepsilon/2$ on $b(\varepsilon),b_{o}]$ for some $\lambda\in(0,1)$ to be specified later. From (1) and Theorem 1 we may conclude that $\hat{\varphi}_{n}(b(\varepsilon))\geq\varphi(b_{o})-\lambda\varepsilon$ und $\hat{\varphi}_{n}\leq\varphi(b_{o})+\lambda\varepsilon$ on $b(\varepsilon),b_{o}]$ with asymptotic probability one. Thus it suffices to show that the event

### v3

@@ Abstract.
It is shown that the nonparametric maximum likelihood estimator of a univariate log-concave probability density satisfies desirable consistency properties in the tail regions. Specifically, let $P$ and $f$ denote the true underlying distribution and density, respectively. If $\hat{f}_{n}$ is the estimated log-concave density, and $\hat{\varphi}_{n}=\log\hat{f}_{n}$, then we specify sequences $(b_{n})_{n\in\mathbb{N}}$ such that $P([b_{n},\infty))\to 0$ at a specific speed, ensuring that the absolute errors or absolute relative errors of $\hat{f}_{n},\ \hat{\varphi}_{n}$ and $\hat{\varphi}_{n}^{\prime}$ converge to zero uniformly on sets $a,b_{n}]$. The main tools, besides characterizations of $\hat{f}_{n}$, are exponential and maximal inequalities for truncated moments of log-concave distributions, which are of independent interest.
@@
Chernov bounds, consistency, exponential inequalities, log-linear densities.
@@ 1 Introduction
Concerning consistency of $\hat{\varphi}_{n}$, let $\{x\in\mathbb{R}:0<F(x)<1\}=:(a_{o},b_{o})$ with $\infty\leq a_{o}<b_{o}\leq\infty$. It was shown by Dümbgen and Rufibach (2009) that for any fixed interval $a,b]\subset(a_{o},b_{o})$, the supremum of $|\hat{\varphi}_{n}-\varphi|$ over $[a,b]$ is of order $O_{\mathrm{p}}(\rho_{n}^{1/3})$, where $\rho_{n}:=\log(n)/n$. If $\varphi$ is Hölder-continuous on a neighborhood of $[a,b]$ with exponent $\beta\in(1,2]$, this rate improves to $O_{\mathrm{p}}(\rho_{n}^{\beta/(2\beta+1)})$. Uniform consistency of $\hat{\varphi}_{n}$ on arbitrary compact subintervals of $(a_{o},b_{o})$ implies that $\int\bigl|\hat{f}_{n}(x)-f(x)\bigr|\,\mathrm{d}x\to_{\mathrm{p}}0$. (Throughout this paper, asymptotic statements refer to $n\to\infty$.) Pointwise limiting distributions at a single point $x_{o}\in(a_{o},b_{o})$ have been derived by Balabdaoui et al. (2009), assuming that $\varphi$ is twice continuously differentiable in a neighborhood of $x_{o}$. Kim and Samworth (2016) showed that the expected squared Hellinger distance between $\hat{f}_{n}$ and $f$ is of order $O(n^{-4/5})$. Numerous further results about $\hat{\varphi}_{n}$ and $\hat{f}_{n}$ have been derived thereafter, including multivariate settings, see the review of Samworth (2018). In the present univariate setting, fast algorithms for the computation of $\hat{\varphi}_{n}$ and related objects are provided by Dümbgen and Rufibach (2011) and Dümbgen et al. (2021). Experiments with simulated data show that even in the tail regions, that is, close to $a_{o}$ and $b_{o}$, the estimator $\hat{\varphi}_{n}$ is surprisingly accurate. In view of this empirical finding, Müller and Rufibach (2009) developed new estimators for extreme value analysis with excellent empirical performance. However, the currently available theory about the asymptotic properties of $\hat{\varphi}_{n}$ does not explain its good performance in the tail regions.
In what follows, several results about $\hat{\varphi}_{n}(x)$ and the right-sided derivative $\hat{\varphi}_{n}^{\prime}(x+)$ for $x$ close to $b_{o}$ are derived. By symmetry, these findings carry over to results in the left tail region. The main results are presented in Section 2 while Sections 3 and 4 provide the proofs. Moreover, the results in Section 3.4 are of independent interest.
@@ 2 Main results
(c) Suppose that $\varphi^{\prime}(\infty-)=$-$\infty$ and $\varphi$ is differentiable on some halfline $(a_{*},\infty)\subset(a_{o},\infty)$ with Lipschitz-continuous derivative $\varphi^{\prime}$. Then, for arbitrary fixed $a\in(a_{*},\infty)$ such that $\varphi^{\prime}(a)<0$,
$\sup_{x\in[a,b_{n}]}\,\Bigl(\frac{\hat{\varphi}_{n}^{\prime}(x+)}{\varphi^{\prime}(x)}-1\Bigl)^{+}\ \to_{\mathrm{p}}\ 0$
and
$\sup_{x\in[a,b_{n}]}\,\bigl(\hat{\varphi}_{n}(x)-\varphi(x)\bigr)^{+}\ \to_{\mathrm{p}}\ 0,\quad\sup_{x\in[a,b_{n}]}\,\frac{\bigl(\varphi(x)-\hat{\varphi}_{n}(x)\bigr)^{+}}{1+|\varphi(x)|}\ \to_{\mathrm{p}}\ 0.$
@@ 3.2 Inequalities for $\hat{P}_{n}^{\mathrm{emp}}$
Note first that for arbitrary indices $0\leq j<k\leq n+1$ with $k-j\leq n$, the random variable $P(X_{(j)},X_{(k)})$ follows a beta distribution with parameters kjk-j and n+1k+jn+1-k+j, see Chapter 3.1 of Shorack and Wellner (1986). Here we set $X_{(0)}:=a_{o}$ and $X_{(n+1)}:=b_{o}$. In particular, its mean equals
@@ 3.3 Truncated and conditional means of $P$
Note that $x_{1}=a$ would imply that $f>\tilde{f}$ on $(a,b)$ or $f<\tilde{f}$ on $(b,\infty)$, and in both cases we would end up with $\int_{a}^{b}f(x)\,\mathrm{d}x\neq\int_{a}^{b}\tilde{f}(x)\,\mathrm{d}x$. Similarly one can exclude the cases $x_{2}=\infty$, $x_{1}\geq b$ and $x_{2}\leq b$. Consequently, we know that there exist constants $a<x_{1}<b<x_{2}<\infty$ such that
@@ 3.4 Exponential and maximal inequalities
If $Q$ follows an exponential distribution, it is well known that the above inequality holds with equality. Now suppose that $Q$ is an arbitrary distribution with log-density $\varphi$. Let $\tilde{\varphi}$ be the log-density of the exponential distribution $\tilde{Q}$ with mean $\mu_{Q}$, that is, $\tilde{\varphi}(x)=-x/\mu_{Q}-\log(\mu_{Q})$ for $x\geq 0$ and $\tilde{\varphi}(x)=$-$\infty$ for $x<0$. Concavity of $\varphi$ and linearity of $\tilde{\varphi}$ on $0,\infty)$ together with equality of the means imply that for suitable real numbers $0<a<b$, $\varphi(x)\leq\tilde{\varphi}(x)$ for $x\not\in[a,b]$ and $\varphi(x)\geq\tilde{\varphi}(x)$ for $x\in(a,b)$. By Lemma b in Karlin and Novikoff (1963) (see also Theorem 3.A.44 in Shaked and Shanthikumar (2007)), we obtain
$\int\Psi\,\mathrm{d}Q\ \leq\ \int\Psi\,\mathrm{d}\tilde{Q}\quad\text{for all convex}\ \Psi:\mathbb{R}\to\mathbb{R}.$ (11)
Applying (11) to $\Psi(x)=\exp(tx/\mu_{Q})$ for arbitrary $t<1$, and using independence of $Y_{1},\ldots,Y_{m}$ concludes the proof. ∎
@@ 4 Proofs of the main results
To this end, we show that for any fixed $\varepsilon>0$, the inequality $\hat{\varphi}_{n}(b_{n})\leq\varphi(b_{o})-\varepsilon$ holds with asymptotic probability zero. Let $b(\varepsilon)\in(a_{o},b_{o})$ such that $|\varphi-\varphi(b_{o})|\leq\lambda\varepsilon/2$ on $b(\varepsilon),b_{o}]$ for some $\lambda\in(0,1)$ to be specified later. From (1) and Theorem 1 we may conclude that $\hat{\varphi}_{n}(b(\varepsilon))\geq\varphi(b_{o})-\lambda\varepsilon$ and $\hat{\varphi}_{n}\leq\varphi(b_{o})+\lambda\varepsilon$ on $b(\varepsilon),b_{o}]$ with asymptotic probability one. Thus it suffices to show that the event